# Principal-Agent Multitasking: the Uniformity of Optimal Contracts and its Efficient Learning via Instrumental Regression

### v1

@@ 5 Improving the Estimation with Repeated Observations and Diversity
Therefore, we can use the following GMM estimator to obtain an estimate $\hat{\theta}$:
@@
The algorithm will simply use the estimator above, without regard to ‘exploration’. In the implementation, the algorithm proceeds in epochs with epochs doubling in length to save computation. In each epoch $e$, the principal sets the contract using the estimation $\hat{\theta}_{e}$ according to the above formula using data from the previous epochs. See Algorithm 2 for details.
@@
At a technical level, the analysis is a combination of the convergence analysis of GMM estimators (e.g. Chapter 18 in [5]) and smoothed analysis in bandit problems ([19, 20, 21]).
@@
Interestingly, the “fast" convergence also holds under a different notion of ‘diversity’. Instead of assuming agents’ cost functions are diverse, consider a setting where agents do not necessarily best respond, and their response $w_{t}$ may be a random vector with some variance lower bounded from below. In this setting, the minimum singular value $\widetilde{X}_{t}^{\top}X_{t}$ still grows linearly with $t$, and the “fast" convergence rate is achieved.
6 Numerical Experiments
Figure 2: (a): Total loss using the Contract as IV, and exploration is achieved through small random perturbations; (b): Total loss using the repeated observations as IV and when the “diversity" condition is satisfied.
This section describes numerical simulations that test the performance of the learning algorithms based on the GMM estimators. In the experiments, the number of tasks is chosen as $d=5$, the total number of rounds is chosen as $T=10^{6}$, and the true parameter $\theta$ is chosen so that each entry is between $[1,3]$. Experiments are ran on a personal laptop.
The experiments have two settings. In the first setting, the contract is used as the instrumental variable, and small random perturbations are added as a means to achieve exploration in the exploration phase (Algorithm 1). The small perturbations are chosen as truncated Gaussian vectors so that the posted contract has components within $[1,3]$. The agents have cost function $c(w)=\norm{w}_{2}^{2}$. The cumulative utility loss of a single run is plotted in Figure 2(a). It can be seen the loss grows quickly in the warm start phase (which consists of $d\sqrt{T}\approx 5\times 10^{3}$ rounds), and then the regret grows slowly in the remaining rounds. The experiments are run for 20 trials, and the average cumulative loss in the 20 trials is 4134.7, roughly on the order of $O(d\sqrt{T})$.
In the second setting, repeated observations are used as the instrumental variable (Algorithm 2). The agent’s cost function is chosen as a diagonal quadratic function with form
$c(w)=\sum_{i=1}^{d}\frac{w_{i}^{2}}{2k_{i}}.$
The parameter $k_{i}$ is sampled from the uniform distribution on $[1,2]$ so that the ‘diversity’ condition is satisfied. The cumulative loss of a single run is plotted in Figure 2(b). The loss grows quickly in the 1st epoch (consisting of roughly $\widetilde{O}(d)$ rounds) and then grows very slowly in the remaining rounds. The experiment is repeated for 20 trials, and the average cumulative loss is 89.6 (recall the theoretical limit is $\widetilde{O}(d)$).
7 Conclusion and Limitation
This work studied the multitask principal-agent problem. I first present a result justifying the use of uniform contracts when the principal contracts with heterogeneous agents. Then, I study the problem of learning the optimal contract and identify the problem as a measurement error problem, and propose using instrumental regression with the generalized method of moments estimator to estimate and learn the optimal contract. A limitation of the current work is that in the “diversity” condition in Section 5, the agents are limited to have quadratic cost functions. I believe this can be extended to more general cost functions (beyond quadratic functions and beyond $k=2$). Specifically, the “diversity” condition will hold whenever there are some variations in the agent’s “abilities" in each task. I will leave a more formal treatment to future work.
Appendix A Proof of Theorem 2
@@
Denote $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and collect $\gamma_{t}$ into the column vector $\Gamma_{e}$.
Proposition 2.
$\norm{\hat{\theta}_{e}-\theta}_{2}\leq\frac{\norm{\widetilde{X}_{e}\Gamma_{e}}}{\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})}$.
@@
With high probability $\delta$, $\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}w_{t}^{\top})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2$.
@@
$\mathbb{E}[W_{t}]$ $=\mathbb{E}[w_{t}w_{t}^{\top}]$
$=\mathbb{E}[\mathtt{diag}(\kappa_{t})\theta_{t}\theta_{t}^{\top}\mathtt{diag}(\kappa_{t})]$
$=\mathbb{E}[\Theta_{t}\kappa_{t}\kappa_{t}^{\top}\Theta_{t}]$
$=\Theta_{t}\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}]\Theta_{t}$
Recall $\lambda_{\min}(\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}])\geq\lambda_{0}$. The following holds
$\lambda_{\min}(\mathbb{E}W_{t})\geq c_{0}^{2}\lambda_{0}.$
@@
$\Pr[\lambda_{\min}(\sum_{k}W_{k})\leq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2]$ $\leq d\cdot\exp(-\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/(8dC_{0}K_{0}))$
@@
|$\absolutevalue{\tau_{e}}\geq\frac{8dC_{0}K_{0}}{c_{0}^{2}\lambda_{0}}\cdot\ln(d/\delta)$
@@
With probability $1-2\delta$, for each epoch $e$ with length as specified in Algorithm 2, $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4$.
@@
In the following denote $\varepsilon_{t}=x_{t}-w_{t},\widetilde{\varepsilon}_{t}=\widetilde{x}_{t}-w_{t}$.
@@
$=\sum_{t\in\tau_{e}}(w_{t}+\varepsilon_{t})(w_{t}+\widetilde{\varepsilon}_{t})^{\top}$
$=\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t}+\varepsilon_{t}^{\top}w_{t}+w_{t}\varepsilon_{t}+\varepsilon_{t}^{\top}\widetilde{\varepsilon}_{t}$
@@
$\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2.$
@@
$\sum_{t\in\tau_{e}}\varepsilon_{t}w_{t}^{\top}+w_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}$
@@
Recall $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and $\Gamma_{e}$ is the column vector collecting the variables $\gamma_{t}$. Note that $\gamma_{t}\widetilde{x}_{t}$ is a zero-mean, $\sigma_{0}C_{0}K_{0}$-subgaussian random vector and independent conditioned on all previous steps. The result then follows from a standard concentration inequality. ∎
@@
By Lemma 4, Lemma 3, and Proposition 2, with probability $1-3\delta$, the regret in each epoch can be upper bounded as |$\absolutevalue{\tau_{e}}\cdot\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2}))=\widetilde{O}(d/\lambda_{0}^{2})$. Choose some appropriate $\delta$, e.g., $\delta<O(1/T^{2})$. Since there are at most $\log T$ epochs, the total regret will be $\widetilde{O}(d/\lambda_{0}^{2})$. ∎

### v2

@@ 7.1 Offline Setting: An Alternative GMM Estimator
Assume the principal is given some offline data as $((\beta_{i},x_{i},\widetilde{x}_{i},y_{i})_{i=1}^{T})$. Then, the principal can use the following GMM estimator to obtain an estimate $\hat{\theta}$:
@@
Here $X_{T},\widetilde{X}_{T}$ are a $T\times d$ matrices where row $t$ is $x_{t}$ or $\widetilde{x}_{t}$ respectively; $Y_{T}$ is a vector with the $t$-th entry being $y_{t}$.
We state the convergence result for the GMM estimator above.
Proposition 4.
With probability $1-\delta$,
$\norm{\theta-\theta^{*}}_{2}\leq\frac{\sqrt{dT\log(dT/\delta)}}{\sigma_{\min}(\widetilde{X}_{T}^{\top}X_{T})}.$
The proof is straightforward and similar to that of Proposition 1.
7.2 Online Setting: Pure exploitation algorithm
We again turn to the online setting where the principal must learn the optimal contract. The algorithm will simply use the estimator above, without regard to ‘exploration’. In the implementation, the algorithm proceeds in epochs with epochs doubling in length to save computation. In each epoch $e$, the principal sets the contract using the estimation $\hat{\theta}_{e}$ according to the above formula using data from the previous epochs. See Algorithm 2 for details.
@@ 7.2 Online Setting: Pure exploitation algorithm
At a technical level, the analysis is a combination of the convergence analysis of GMM estimators (e.g. Chapter 18 in [Greene, 2008]) and smoothed analysis in bandit problems ([Bastani et al., 2021, Kannan et al., 2018, Sivakumar et al., 2022]).
8 Proof of Robustness Result in Theorem 1
We will show that for any contract $w$, there exists an non-negative affine contract $\phi$ and compatible distributions $\mathcal{F}$ such that
$u(\phi;\mathcal{F})\geq u(w;\mathcal{F}).$
Since under an affine contract, the agent’s expected payment for any action $a$ remains the same for any distribution, we know
$u(\phi)=u(\phi;\mathcal{F}).$
Moreover, by definition
$u(w)\leq u(w;\mathcal{F}).$
By the above we know that
$u(\beta)\geq u(w).$
Then using Lemma 1, which shows we can further improve an affine contract to a linear contract, the theorem is proven.
We define the following set of points:
[…]
For each upper hyperplane $H$ supporting the set $S$ from above, consider the set of points that the plane $H$ supports, and project these points back to the space $\mathcal{X}$. Denote this set as $P(H)$. By definition $P(H)\subset\mathcal{X}$.
By a classic result in convex geometry (e.g. [Lee and Santos, 2017], [De Loera et al., 2010]), if we consider the set of upper facets of the set $S$ (for which each defines an upper hyperplane), projecting each facet back to $\mathcal{X}$ in fact gives an exact partition of $\mathcal{X}$, i.e.,
$\cup_{H:H\text{ touches an upper facet}}\mathtt{conv}(P(H))=\mathtt{conv}(\mathcal{X}).$
For our purposes, we will be considering upper supporting hyperplanes whose normal vector’s last component is bounded away from 0. lower bounded by some $\varepsilon$. In the following denote $n_{d+1}(H)$ to as the last component of the normal vector. For hyperplanes with $n_{d+1}(H)=0$, they do not represent valid contracts.
Lemma 2.
[…]
$\cup_{H\in\mathcal{H}}\mathtt{conv}(P(H))=\mathtt{conv}(\mathcal{X}).$
Proof.
The set $S$ has at most a finite number of upper facets, to which each corresponds to a hyperplane $H$. Then, simply set $\varepsilon$ to be the minimum of $n_{d+1}(H)$ among all such $H$, and the lemma is proven. ∎
Now, it can be seen that the set $\mathcal{H}$ as in the previous lemma is topologically equivalent to a closed and convex set. This will allow us to apply fixed-point theorems in the subsequent analysis.
In the following, we define $\phi(H)$ to be the affine contract associated with the hyperplane $H$, and use $a(H)=a(\phi(H))$ to denote the action that is induced by $\phi(H)$.
Lemma 3.
There exists an upper supporting hyperplane $H$, such that the incentivized action $a^{*}$ under $\phi(H)$ lies in the set $\mathtt{conv}(P(H))$.
Recall this is the notion of “self-owned” hyperplanes as introduced previously.
Proof.
By the previous lemma, the union of all $\mathtt{conv}(P(H))$ cover the entire space $\mathcal{A}$. Then for any $a$, there must exist some hyperplane $H^{\prime}$, such that $a(H)\in\mathtt{conv}(P(H^{\prime}))$.
So, let us define a function $g:\mathcal{H}\rightarrow 2^{\mathcal{H}}$, such that for each upper supporting hyperplane $H$, the function maps to a set of supporting hyperplanes defined as follows:
[…]
By the previous lemma, the set $g(H)$ is non-empty. Further, we can see that the set $g(H)$ is convex and that $g$ has a closed graph (see Section A.3 for a brief discussion). By Katukani’s fixed point theorem, there exists some hyperplane $H$ such that $g(H)=H$. Hence the theorem is proven. ∎
We can now show that there exists some affine contract which does no worse than $w$ in worse-case.
Lemma 4.
Let $H$ be the hyperplane as in the statement of the previous lemma (i.e., “self-owned”) and $H$ induces $a(H)$. Then there exists some distribution under which the contract $w$ also induces $a(H)$. Further, $\phi(H)$ uses no greater payment than $w(\cdot)$.
Proof.
First note that under any distribution, for any action $a\in\mathcal{A}$, the payment to the agent under $w$ is certainly no larger than under $\phi$, since $\phi$ bounds $w$ from above. Now, since $H$ is “self-owned”, we have $a(H)\in\mathtt{conv}(P(H))$. We can define a compatible distribution where the signal distribution of action $a(H)$ is supported only on points in $P(H)$. For points $x\in P(H)$ we have $\phi(x)=w(x)$. Then, the payment to the agent if he takes action $a(H)$ is exactly the same under $\phi(H)$ as under $w(\cdot)$. Hence, under this distribution, the contract $w$ also induces action $a(\beta)$. ∎
Thus so far, we have shown that any contract $w(\cdot)$ can be improved to an affine contract whose worst-case payoff is no less than that of $w(\cdot)$. By Lemma 1, any affine contract can be improved to a linear contract. This completes the proof.
Our main result in this section Theorem 1 then follows from the above lemmas.
@@ 8 Proof of Robustness Result in Theorem 1
Our results can also be extended to the case when the signal space is infinite, e.g., $\mathcal{X}=[0,1]^{d}$. However, this requires some additional regularity assumptions. In particular, we need to ensure the contract does not grow “too quickly” at the boundary. This avoids the case that at the boundary a “vertical” supporting hyperplane occurs, i.e., one with $n_{d+1}(H)=0$. Such hyperplanes does not correspond to valid contracts, and Lemma 2 does not hold.
9 Proof of Regret in Theorem 3
@@
Denote $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and collect $\gamma_{t}$ into the column vector $\Gamma_{e}$. Recall that by Proposition 4, we have
$\norm{\hat{\theta}_{e}-\theta}_{2}\leq\frac{\norm{\widetilde{X}_{e}\Gamma_{e}}}{\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})}.$ (2)
@@
With high probability $\delta$, $\lambda_{\min}(\sum_{t\in\tau_{e}}a_{t}a_{t}^{\top})\geq\absolutevalue{\tau_{e}}\lambda_{0}/2$.
@@
By the diversity condition,
$\lambda_{\min}(\mathbb{E}W_{t})\geq\lambda_{0}.$
@@
$\Pr[\lambda_{\min}(\sum_{k}W_{k})\leq\absolutevalue{\tau_{e}}\lambda_{0}/2]$ $\leq d\cdot\exp(-\absolutevalue{\tau_{e}}\lambda_{0}/(8K_{0}))$
@@
|$\absolutevalue{\tau_{e}}\geq\frac{8K_{0}}{\lambda_{0}}\cdot\ln(d/\delta)$
@@
With probability $1-2\delta$, for each epoch $e$ with length as specified in Algorithm 2, $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})\geq\absolutevalue{\tau_{e}}\lambda_{0}/4$.
@@
In the following denote $\varepsilon_{t}=x_{t}-a_{t},\widetilde{\varepsilon}_{t}=\widetilde{x}_{t}-a_{t}$.
@@
$=\sum_{t\in\tau_{e}}(a_{t}+\varepsilon_{t})(a_{t}+\widetilde{\varepsilon}_{t})^{\top}$
$=\sum_{t\in\tau_{e}}a_{t}^{\top}a_{t}+\varepsilon_{t}^{\top}a_{t}+a_{t}\varepsilon_{t}+\varepsilon_{t}^{\top}\widetilde{\varepsilon}_{t}$
@@
$\lambda_{\min}(\sum_{t\in\tau_{e}}a_{t}^{\top}a_{t})\geq\absolutevalue{\tau_{e}}\lambda_{0}/2.$
@@
$\sum_{t\in\tau_{e}}\varepsilon_{t}a_{t}^{\top}+a_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}$
@@
Recall $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and $\Gamma_{e}$ is the column vector collecting the variables $\gamma_{t}$. Note that $\gamma_{t}\widetilde{x}_{t}$ is a zero-mean, $\sigma_{0}K_{0}$-subgaussian random vector and independent conditioned on all previous steps. The result then follows from the standard concentration inequality (see Section A.1). ∎
@@
By Lemma 8, Lemma 7, and Equation 2, with probability $1-3\delta$, the regret in each epoch can be upper bounded as |$\absolutevalue{\tau_{e}}\cdot\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2}))=\widetilde{O}(d/\lambda_{0}^{2})$. Choose some appropriate $\delta$, e.g., $\delta<O(1/T^{2})$. Since there are at most $\log T$ epochs, the total regret will be $\widetilde{O}(d/\lambda_{0}^{2})$. ∎
10 Conclusion
In this work, we examined the role of linear contracts in principal-agent problems from three key perspectives: robustness, uniformity, and learning.
Firstly, we analyzed the robustness properties of linear contracts. In an ambiguous principal-agent setting where only first-moment information is available, we demonstrated that a linear contract is worst-case optimal. This result underscores the resilience of linear contracts under uncertainty, providing a rationale for their widespread use in practical applications. A natural question for future research is to investigate the extent to which our assumptions can be relaxed while preserving the robustness guarantees. Understanding the boundary conditions of these results could yield deeper insights into the trade-offs between contract simplicity and performance in uncertain environments.
Second, we provided a theoretical justification for the use of uniform contracts when contracting with heterogeneous agents. Our analysis establishes conditions under which uniform contracts are not only feasible but also optimal. Specifically, when agents exhibit certain cost function properties, a uniform contract structure can align incentives effectively without the need for agent-specific tailoring. This result is particularly relevant in settings such as franchising or platform-based economies, where principals interact with diverse agents yet often employ standardized contracts.
Finally, we investigated the problem of learning optimal contracts in scenarios where the principal lacks complete knowledge of agent-specific parameters. We drew a connection between contract learning and instrumental regression, highlighting how techniques from causal inference can inform strategic contract design. Notably, recent research has recognized the relationship between learning in strategic settings and causal inference methods, such as instrumental regression. Building on this, a promising direction for future work is to explore how learning with strategic interactions in Stackelberg games can be framed within the instrumental regression framework. This could open new avenues for leveraging econometric techniques in the study of contract optimization.
Appendix A Appendix
A.1 Concentration Lemma
Lemma 9.
Let $\gamma_{1},\dots,\gamma_{T}$ be independent $\sigma$-subgaussian random variables. Let $v_{1},\dots,v_{T}$ be vectors in $\mathbb{R}^{d}$ with each $v_{T}$ chosen arbitrarily as a function of $(v_{1},\gamma_{1}),\dots,(v_{t-1},\gamma_{t-1})$ subject to $\norm{v_{t}}\leq C$. Then with probability at least $1-\delta$,
$\norm{\sum_{t=1}^{T}\gamma_{t}v_{t}}\leq\sqrt{2dC\sigma T\log(Td/\delta)}.$
A.2 Proof for Proposition 2
Assumption 8.
There exists some constant $\underline{a}>0$, such that under the contract $\beta=e_{i}$, any agent’s private effort satisfies $a_{i}>\underline{a}$.
Proof of Proposition 2.
Fix a dimension (task) $i\in[d]$. Consider the $O(\sqrt{T})$ rounds in which the principal posted the contract $\beta=\mathbf{e}_{i}$. Let the sum of the signal for the $i$-th task be $X_{i}$ over these rounds, and let the sum of agents’ true effort be $A_{i}$.
Note that in the denominator of the GMM estimator, the matrix $B_{\tau}X_{\tau}$ simply becomes $\mathtt{diag}(X_{1},\dots,X_{d})$. Using proposition 1, we have with probability $1-\delta$,
$\norm{\hat{\theta}-\theta}_{2}\leq\frac{\sqrt{d\sqrt{T}\log(1/\delta)}}{\min(X_{i})}.$
Hence let us now turn to bound $\min(X_{i})$. Note that with probability $1-\delta_{0}$, by a standard Hoeffding inequality, we have
$\Pr[\absolutevalue{X_{i}-W_{i}}<O\left(\sqrt{\frac{\sigma^{2}\ln(1/\delta_{0})}{T^{1/2}}}\right)]<\delta_{0}.$
Hence choosing $\delta_{0}=\delta/d$, with probability at least $1-\delta$, for large enough $T$, we have $\min_{i}X_{i}\geq\Omega(T^{1/2})$.
In the initial exploration rounds, the estimation error after the initial $d\sqrt{T}$ rounds is:
$\norm{\hat{\theta}-\theta^{*}}_{2}=O(\sqrt{d\log(1/\delta)/\sqrt{T}}).$
The regret in subsequent rounds can be bounded as:
$T\cdot\norm{\hat{\theta}-\theta^{*}}_{2}^{2}=O(d\sqrt{T}\cdot\log(1/\delta)).$
Hence the total regret is $O(d\sqrt{T}\cdot\log(1/\delta))$. ∎
A.3 Application of Kakutani’s Fixed-Point Theorem
Applying Kakutani’s fixed-point theorem requires the following conditions on the mapping $g:\mathcal{H}\rightarrow 2^{\mathcal{H}}$. For any $x$, $g(H)$ is non-empty, convex, and closed. Further $\phi$ has is upper hemicontinuous (which together with closeness of $g(H)$ implies $g$ have a close graph).
One of the key conditions, non-emptiness, is guaranteed by Lemma 2. To see that $g(H)$ is convex, note that if two hyperplanes $H_{1}$ and $H_{2}$ supports $a(H)$, so does their convex combination. To see $g(H)$ is closed, note that for any sequence of hyperplanes $(H^{\prime}_{i})$ which converge to $H^{\prime}$, if all hyperplanes in the sequence $H^{\prime}_{i}$ supports $a(H)$, then so does the limit $H^{\prime}$.
To show upper hemicontinuity, let $(H^{\prime}_{i})$ be a sequence of hyperplanes that converge to $H^{\prime}$, and let $(H_{i})$ be a sequence of points that converge to $H$. Note that if for any $i$ we have that $H^{\prime}_{i}$ supports $a(H_{i})$, then in the limit $H^{\prime}$ supports $a(H)$. Hence upper hemicontinuity is shown.